# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v3

@@ 4.2.5 Multi-Modal Representations
5 Case Studies of TDA for NLP Tasks
We showcase the utility of TDA in non-trivial NLP tasks by applying persistent homology and mapper using the Theoretical and Non-theoretical approaches. See below:
5.1 Theoretical Approaches.
(Semantic) Topological Space.
Although applications of this framework have yielded interesting findings, which are discussed in Section 3.1.1, it might still be interesting to answer the question - does topological features of text capture semantic features as well? Thus, we perform a small analysis to illustrate a real-world application. We perform this analysis on a relevant topic - distinguishing between watermarked and non-watermarked LLM-generated texts.
We extract topological and linguistic features for a correlation test. The text is initially represented numerically with BERT’s attention embeddings, then as an undirected graph, similar to Kushnareva et al. (2021) to extract topological features. Next, using the undirected graph as input, we calculate topological features in the 0-dimension, such as the number of non-zero values, maximum value, mean of values, number of values, and persistent entropy. Also, for the linguistic features, we extract entropy, perplexity, position of noun & adjective, coherence, flesh reading ease score & grade, number of stopwords, length of text, and number of unique words. Finally, we use distance correlation to measure the statistical dependence of the two non-linear sets of features. See Figures 6(a) & 6(b) for the correlation matrix heatmap. Many of the TDA features correlate with the linguistic features, with the highest correlation being the length of text (doc_length) and number of topological values in $\beta_{0}$ (betti_curve_0). Lastly, these results suggest that topological features can capture semantic features, however, a more comprehensive study is needed to make this a strong finding. In addition, such a finding can inform the non-theoretical approaches we adopt for real-world NLP tasks.
Topological “Shape” of Words.
Using the same method as above to represent the Watermarked vs. Non-Watermarked texts, we visualize the “shape” of their data using Mapper to observe the similarity or distinctness between the text types. See Figures 7(a) and 7(b) for the Mapper plots of the Watermarked and Non-Watermarked texts, respectively. We observe from the Figures that these different texts are very similar, although the watermarked text has slightly more nodes and edges than the non-watermarked texts, suggesting more connectivity. Thus, we can conclude that these texts are topologically similar, which could suggest that the watermarking technique does not significantly deviate from the style and quality of the non-watermarked version. However, to further explain these results, a more comprehensive study is required to interpret topological features extracted from text linguistically. This further stresses the validity of the theoretical study in Section 5.1. In addition, such findings could indicate that the watermarking technique does not break the semantics, style, and syntax of the non-watermarked pre-trained model.
@@ 5.2 Non-Theoretical Approaches
Model Interpretation and Analysis.
To illustrate an interesting task for applying TDA to NLP, we aim to investigate whether the embeddings of English and non-English text pairs are the same. This analysis can provide insights into the ability of text embeddings to capture the same information but in different languages. The idea is to represent an English text with a monolingual English embedding model and a non-English text with the same meaning with a multilingual embedding model, and compare the similarities between these embeddings. This similarity is captured by calculating both the Euclidean and topological distances. For the topological distance, we use Wasserstein distance, a similarity metric between two probability distributions, typically used in computational topology Panaretos and Zemel (2019); Shin (2019); Draganov and Skiena (2024). Specifically, we evaluate the Euclidean and Wasserstein distances between the embeddings of English-French translation text pairs. For the Euclidean distance, we compare the embeddings from the English model and the multilingual model. To calculate the Wasserstein distance, we first extract topological features from the embeddings, represent them as persistent diagrams, and then calculate the distance. Euclidean and Wasserstein distance range is [0, $\infty$). We further normalize the embeddings such that both distance metrics are comparable. See Table 2 for the distance values. Thus, we obtain an average distance of 1.412 and 0.054 for the Euclidean and Wasserstein distances, respectively. These values suggest that these embeddings capture the similarity between the text pairs, with Wasserstein being almost zero, meaning that the text pair embeddings are topologically similar. Finally, we propose a new way of evaluating the strength of multilingual models by observing their topological similarity to monolingual models.
@@ 6 Open Problems and Future Directions
6.1 Pros and Cons of TDA in NLP
There are several benefits and challenges to adopting TDA techniques to solve non-trivial problems. We outline these pros and cons for NLP problems in Table 3. The most glaring benefit and challenge are that TDA is effective in low-resource settings and the high computational cost of employing TDA, respectively. This benefit makes TDA particularly attractive for problems with noisy, insufficient, heterogeneous, and high-dimensional datasets. Although the high computational cost presents challenges for the adoption of TDA, recent improvements in computational resources mitigate this challenge. Finally, below we discuss how researchers can leverage these benefits to address the open problems and future directions.
@@ 6.2 LLMs to Convert TDA Concepts to Codes
One of the major challenges in applying TDA to NLP tasks is the steep learning curve associated with its mathematical foundations, which are often accessible only to expert audiences. Moreover, the theorists who develop and understand these advanced concepts do not always collaborate with computational scientists to translate them into executable code. To address this gap, Liu et al. (2024) proposes leveraging ChatGPT to generate Python code for TDA concepts by training it on these mathematical foundations. Their findings suggest that ChatGPT can alleviate this bottleneck, particularly for complex TDA concepts like hypergraphs, digraphs, and persistent harmonic space, which have not been as heavily explored as the Vietoris-Rips complex Liu et al. (2024). Similarly, experts can develop specialized code generators, such as fine-tuning models like Code-Llama Roziere et al. (2023) on TDA concepts. Creating a dedicated LLM for TDA code generation could significantly lower the entry barrier, encouraging the NLP community to explore TDA applications more innovatively.
@@ 6.3 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigate the shape of words and their embeddings in Indo-European languages and find similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts as in languages clustered closely together are similar or influenced by the other. Finally, we observe only 12 theoretical TDA works in NLP, compared to over 80 non-theoretical ones. Thus, we need more theoretical TDA approaches as the depth and understanding of performance from a topological perspective without more investigations.
@@ 6.4 TDA for Interpretability
Interpreting TDA features in NLP problems given its non-intuitive nature is very challenging. This is evident in the fact that most TDA for explainability applications is mostly in Computer Vision Saul and Arendt (2018), where the structure is distinct. Consequently, the interpretation of TDA features for text or speech data remains an open problem. There are currently two main tasks in this space - (1) explain model performance by interpreting TDA features; (2) explain model performance by using TDA to probe the prediction space or data. Either tasks require a deeper understanding of TDA such that intuitive explanations can be used to tie topology to linguistic phenomena. Specifically, we need novel approaches that link TDA features to linguistic phenomena, for instance, disentangling $\beta_{0}$, and $\beta_{1}$’s representations to different properties of natural texts such as coherency, and writing style. This can be done either through visualizing the prediction space of a model Rathore et al. (2023) or probing the prediction space of models with TDA Gardinazzi et al. (2024); Xenopoulos et al. (2022); Solunke et al. (2024).
@@ 6.5 Novel Applications of TDA
When we have more theoretical approaches of TDA and issues barring the application of TDA on interpretable NLP tasks are mitigated, we can hope that TDA can be applied to even more novel, diverse and important tasks. From Section 4.1, we can see that TDA has been applied to 7 categories on non-theoretical NLP tasks. While many of the tasks are interesting, especially the speech & music processing and health applications, there are still nuanced niche fields that could benefit from TDA. One glaring application is on multi-lingual tasks, just as Haim Meirom and Bobrowski (2022) and García-Castellanos et al. (2024). Due to the benefits of TDA which include performing robustly on heterogeneous, imbalanced, and noisy data, its application on multi-lingual tasks is necessary. Other applications include: semantic and syntactic structure, forensic authorship, embedding space, and multi-modal (e.g., language-vision model) analysis
@@ 6.7 Adversarial Robustness of TDA Features
Robustness to noise, particularly adversarial perturbations, has been an important research topic in NLP. While such robustness of TDA features is promising, there have been only a few works in this direction Perez and Reinauer (2022); Chauhan and Kaul (2022). For instance,  Perez and Reinauer (2022) show that their topologically-augmented BERT model is much more robust than the base BERT model when tested against perturbations generated by TextAttack Morris et al. (2020).  Chauhan and Kaul (2022) also shows that there are some weak correlations between persistent homology features of a trained BERT model and its adversarial robustness against several state-of-the-art attackers. However, all existing works only evaluate on BERT model with simple attack mechanism, missing other security scenarios such as poisoned or backdoor attacks. Therefore, we call for a more comprehensive robustness evaluation of TDA on NLP models, especially from the NLP, ML, and security communities.
@@ 6.8 Topological Deep Learning for NLP
Due to the benefits of TDA and deep learning, a new niche field is born - Topological Deep Learning (TDL) a “the collection of ideas and methods related to the use of topological concepts in deep learning” Papamarkou et al. (2024). TDL allows a deep learning model to be integrated more deeply with concepts of algebraic topology, such as the introduction of simplicial neural networks (NNs), which are NNs with layers made up of simplicial complexes. This deeper integration of TDA into NNs makes TDL particularly useful for the explosion of high-dimensional data. These high-dimensional data need better tools for processing as the current tools shrink the dimension, resulting in information loss. In NLP, one particular approach to integrate TDA with high-dimensional NLP embeddings has been the utilization of text in graphical forms, which have been shown to yield better results than directly using texts as a sequence of tokens Zhong et al. (2020); Liu et al. (2023); Phan et al. (2023). Nevertheless, more research is still needed to validate such an approach.

### v4

@@ 6.2 Non-Theoretical Approaches
To illustrate an interesting task for applying TDA to NLP, we aim to investigate whether the embeddings of English and non-English text pairs possess similar topological structures. This analysis can provide insights into the ability of text embeddings to capture similar structural information across different languages. The idea is to represent an English text with a monolingual English embedding model and a non-English text with the same meaning as a multilingual embedding model, and compare the similarities between these embeddings. This similarity is captured by calculating both the Euclidean and topological distances. For the topological distance, we use Wasserstein distance, a similarity metric between two probability distributions, typically used in computational topology (Panaretos & Zemel, 2019; Shin, 2019; Draganov & Skiena, 2024). Specifically, we evaluate the Euclidean and Wasserstein distances between the embeddings of English-French translation text pairs [^8] https://huggingface.co/datasets/aircrypto/English-French-Translations (using the first 500 text pairs). For the Euclidean distance, we compare the embeddings from the English model and the multilingual model using the popular Sentence Transformer family encoders [^9] https://huggingface.co/sentence-transformers/. To calculate the Wasserstein distance, we first extract topological features from the embeddings, represent them as persistent diagrams, and then calculate the distance. Euclidean and Wasserstein distance range is [0, $\infty$). We further normalize the embeddings such that both distance metrics are comparable.
Table 4 shows an average distance of 1.412 and 0.054 for the Euclidean and Wasserstein distances, respectively, both of which are relatively low values. These values suggest that these embeddings are able to capture the similarity between the text pairs, with Wasserstein being almost zero, demonstrating that the text pair embeddings are topologically similar, although one is multilingual and one is monolingual. This also suggests that TDA techniques can be used to evaluate the strength of multilingual models by observing their topological similarity to each individual monolingual embeddings.
@@
7 When to use TDA in NLP
Quick decision rules
  • Try TDA when multi-scale shape or connectivity of an embedding point-cloud (neighborhoods, loops, components) plausibly encodes a signal that the usual similarity pipelines miss (e.g., polysemy or author classification).
  • Do not reach for TDA as a first pass for large end-to-end classification problems where massive supervised models already dominate unless you plan to use TDA as an orthogonal feature, regularizer, or diagnostic; gains are typically complementary, not a wholesale replacement.
  • Use TDA analysis for explainability (what shape does an embedding produce?) or as a targeted augmentation (topological summaries + standard features), not as a black-box substitute for representation learning.
Contexts where TDA is likely to help
  • Word-sense, polysemy detection, or sense induction: Local neighborhood topology around a word can reflect multiple senses (multiple connected components). Jakubowski et al. (2020) show topological measures of polysemy correlate with sense counts and produce competitive sense-induction results.
  • Authorship or stylistic signature (small-to-moderate corpora): Authors may induce distinct global connectivity patterns across entity- or function-word clouds; persistent homology has been used as document-level topological signatures for novelist classification.
  • Storyline, discourse structure or topic evolution: Mapping sentences or paragraphs into embeddings and tracking multi-scale connectivity can reveal hierarchical or story structure (e.g., “story trees”).
  • Data augmentation or generative-data assessment: TDA can quantify whether augmented/generated examples occupy the same topological region as original data (useful for judging augmentation quality). Recent work uses TDA to compare LLM-augmented vs. embedding-similarity augmentation.
  • Robustness, anomaly, or singularity detection in embeddings: Topology-based measures (e.g., persistent H0, H1 summaries) can detect singularities or anomalous neighborhoods (useful for debugging or filtering embeddings).
  • Representation regularization or topology-preserving learning: Adding topological losses (or using topological autoencoders) helps preserve connectivity and global shape in learned latent spaces — helpful when preserving manifold structure matters.
Contexts where TDA is unlikely to help
  • Very large-scale classification where you can train huge supervised models and compute resources are limited
  • Problems where pairwise similarity or local clustering fully captures the signal
  • When you must extract features in real-time on resource-limited hardware
8 Discussion
8.1 Pros and Cons of TDA Applications in NLP
There are several benefits and challenges in adopting TDA techniques to solve non-trivial problems. We outline these pros and cons for various NLP problems in Table 5. The most glaring benefit and challenge are that TDA is effective in low-resource settings and the high computational cost associated with employing TDA, respectively. This benefit makes TDA particularly attractive for problems with noisy, insufficient, heterogeneous, and high-dimensional datasets. Beyond this, TDA can capture both global and local text structures, uncover complex relationships, and is robust to noise. However, issues such as interpretability, limited software support, and the need for specialized expertise still limit its widespread adoption in NLP. Although the high computational cost presents challenges for the adoption of TDA, recent advancement and improved accessibility of computational resources provide some relief. Section 9 will discuss how researchers can leverage these benefits to address the open problems and future directions.
8.2 Other non-TDA Topological Approaches Applied in NLP
There are other topological approaches for extracting structural information from data. Similar to TDA, these approaches also borrow from the mathematical field of algebraic topology, utilizing some of TDA’s concepts including simplicial homology and Morse theory. However, they are not as rigorous as Persistent Homology.
Previous works have utilized such techniques in assessing document coherency by using the connected component dimension ($\beta_{0}$), representing the semantic topological space (Chiang, 2007). Similarly, Santacana (2025) uses a similar framework to build a framework for topological dialogue semantics in the Wolfram Language (i.e., a high-level, symbolic programming language). Next, Ionescu et al. (2025) proposes a framework known as generative topolinguistics, which projects texts to a topological semantic space to understand sociolinguistic phenomena in LLMs. Additionally, to extract the shape of text, Luong et al. (2007), considers text as a topological space by capturing the topology of linguistic features in texts. Which is used to identify the important stylistic features unique to specific authors or genres, specifically the frequent verbs that can be found in literature Luong et al. (2007).
From these applications, it is evident that all applications adopt the theoretical approaches used to reveal and confirm linguistic phenomena. This suggest the importance of topological methods in probing language to find confirmation of known historical linguistic knowledge and insights. Techniques such as this are useful on novel tasks such as discovering the similarities between high-resource languages and low-resource languages, which could potentially help save some of these dying languages.
8.3 Insights into Interesting Applications of TDA in NLP
Theoretical Approaches.
We observe that the most popular application of TDA in NLP has been confirmation of the topologically relationship between language families. For instance, Port et al. (2018; 2022) using the framework of a syntactic topological space and Draganov & Skiena (2024); Dong (2024), topological “shape” of words, they observe and confirm relationships between different language families. In the syntactic topological space, the aim is to capture the syntactic structure of language families from a topological perspective, such that languages in the same family have similar structures. For example, Port et al. (2022) finds that the syntactic topological structure of the Indo-European language family confirms the historical linguistic phenomena that the Hellenic branch played a role in its development. Next, for the topological “shape” of words, the aim is to capture the topological shape of these languages, such that languages of the same family have similar shapes. Moreover, Draganov & Skiena (2024)’s topological shapes of Indo-European languages confirm Port et al. (2018)’s findings.
Thus, by applying TDA to such tasks, we may find new linguistic phenomena which can inform how we numerically represent texts in the future, such that it captures all relevant linguistic principles.
Non-Theoretical Approaches.
We suggest that the most interesting application of the non-theoretical approaches of TDA in NLP is in the model interpretation & analysis task. Researchers particularly utilize TDA features extracted from numerical representations of texts (i.e., mostly Transformer-based embeddings) to analyze LLM outputs in order to make them less opaque. See the subtasks explored under the model interpretation and analysis task:
  1. 1.
    Comparing text representations & embeddings (Haim Meirom & Bobrowski, 2022)
  2. 2.
    Explaining model performance (Spannaus et al., 2024)
  3. 3.
    Testing the manifestation of intelligence and understanding in models (Zadrozny, 2021a)
  4. 4.
    Dynamic weighting for building ensemble models (Proskura & Zaytsev, 2024)
  5. 5.
    Probing LLMs to estimate class separability of text datasets (Gourgoulias et al., 2024)
  6. 6.
    Visualizing the training process of transformer-based models (Rathore et al., 2023)
  7. 7.
    Model compression by pruning redundant layers (Gardinazzi et al., 2024; Balderas et al., 2025)
  8. 8.
    Probing the correlation between sentence vectors and their semantics (Sun & Nelson, 2023)
  9. 9.
    Capturing the homology of hidden representations of BERT (Chauhan & Kaul, 2022)
  10. 10.
    Zero-shot model stitching (i.e., Topological densification) (García-Castellanos et al., 2024)
  11. 11.
    Investigating the topological structure of how ChatGPT makes fair and unfair decisions (Fitz, 2023)
  12. 12.
    Probing BERT’s hidden weights on new grammatical features (i.e., Polypersonality) (Kudriashov et al., 2024)
  13. 13.
    Uncertainty estimation (Kostenok et al., 2023)
  14. 14.
    Comparing the alignment of text and speech embeddings (Yessenbayev & Kozhirbayev, 2022; 2024)
  15. 15.
    Investigating the latent space representation of adversarial vs. non-adversarial texts in LLMs (Fay et al., 2025)
  16. 16.
    Investigating the geometry of textual data augmentation (Feng et al., 2024)
These subtasks illustrate novel ways in which researchers use TDA to probe the hidden weights of AI models in order to make them less opaque. This suggests that while TDA features are not intuitively explainable, it can be used to probe model weights such that model performance is interpretable.
8.4 Differences between standard NLP techniques and TDA
8.4.1 Qualitative differences
Sensitivity to global structure:
Standard embeddings (i.e., Word2Vec, BERT) primarily encode local co-occurrence or contextual similarity. TDA (via persistent homology) captures global and multi-scale structure — e.g., loops, voids, or connected components in the embedding space. For instance, in semantic drift analysis, TDA can detect loops corresponding to cyclical usage of terms over time (i.e., “cloud” in meteorology vs. computing), which embeddings alone treat as proximity without structure.
Topological invariants vs. vector similarity:
Embeddings compare objects with cosine similarity or Euclidean distance. TDA computes persistence diagrams/barcodes that remain invariant under continuous deformations (scaling, rotation). For instance, clusters of synonyms may look similar in embedding space, but TDA can reveal whether the cluster is contractible (one component) or contains holes reflecting polysemy.
8.4.2 Quantitative differences
Dimensionality reduction robustness:
In experiments, persistent homology applied to embeddings yields stable Betti curves that remain informative even after strong dimensionality reduction — unlike raw embeddings, which often degrade in discriminative power.
Additional discriminative signal:
Studies (Hofer et al., 2020) show that adding persistence summaries to standard embedding features improves NLP classification tasks (sentiment, authorship). This suggests that TDA captures information orthogonal to embeddings — otherwise, no performance gain would occur.
Polysemy detection:
Embeddings often “average out” multiple senses of a word. Persistent homology can quantify the multiplicity of connected components or loops in local neighborhoods, providing measurable evidence of polysemy (Jakubowski et al., 2020).
@@ 9 Open Problems and Future Directions
We discuss the open problems and future directions for TDA applications in NLP, as well as ways in which researchers can leverage these benefits and mitigate these risks.
@@ 9.1 LLMs to Convert TDA Concepts to Codes
One of the major challenges in applying TDA to NLP tasks is the steep learning curve associated with its mathematical foundations, which are often accessible only to expert audiences. Moreover, theorists who develop and understand these advanced concepts do not always collaborate with computational scientists to translate them into executable code. To address this gap, Liu et al. (2024) proposes leveraging ChatGPT to generate Python code for TDA concepts by training it on these mathematical foundations. Their findings suggest that ChatGPT can alleviate this bottleneck, particularly for complex TDA concepts like hypergraphs, digraphs, and persistent harmonic space, which have not been as heavily explored as the Vietoris-Rips complex (Liu et al., 2024). Similarly, experts can develop specialized code generators, such as fine-tuning models like Code-Llama (Roziere et al., 2023) on TDA concepts. Creating a dedicated LLM for TDA code generation could significantly lower the entry barrier, encouraging the NLP community to explore TDA applications more innovatively.
@@ 9.2 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov & Skiena (2024) investigates the shape of words and their embeddings in Indo-European languages and finds similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They find that TDA features represent historical facts, such that languages clustered closely together are similar or influenced by each other. Finally, we observe only 11 theoretical TDA works in NLP, compared to over 80 non-theoretical ones. Thus, we need more theoretical TDA approaches, as the depth and understanding of performance from a topological perspective, without further investigation, requires interpretation.
@@ 9.3 TDA for Interpretability
Interpreting TDA features in NLP problems, given its non-intuitive nature is very challenging. This is evident in the fact that most TDA for explainability applications is mostly in Computer Vision (Saul & Arendt, 2018), where the structure is distinct. Consequently, the interpretation of TDA features for text or speech data remains an open problem. There are currently two main tasks in this space - (1) explain model performance by interpreting TDA features; (2) explain model performance by using TDA to probe the prediction space or data. Either tasks require a deeper understanding of TDA such that intuitive explanations can be used to tie topology to linguistic phenomena. Specifically, we need novel approaches that link TDA features to linguistic phenomena, for instance, disentangling $\beta_{0}$, and $\beta_{1}$’s representations to different properties of natural texts such as coherency, and writing style. This can be done either through visualizing the prediction space of a model (Rathore et al., 2023) or probing the prediction space of models with TDA (Gardinazzi et al., 2024; Xenopoulos et al., 2022; Solunke et al., 2024).
@@ 9.4 Novel Applications of TDA
When we have more theoretical approaches of TDA and issues barring the application of TDA on interpretable NLP tasks are mitigated, we can hope that TDA can be applied to even more novel, diverse and important tasks. From Section 5.1, we can see that TDA has been applied to 7 categories on non-theoretical NLP tasks. While many of the tasks are interesting, especially the speech & music processing and health applications, there are still nuanced niche fields that could benefit from TDA. One glaring application is on multi-lingual tasks, just as Haim Meirom & Bobrowski (2022) and García-Castellanos et al. (2024). Due to the benefits of TDA which include performing robustly on heterogeneous, imbalanced, and noisy data, its application on multi-lingual tasks is necessary. Other applications include: Topology-aware neural networks, Topological interpretability, semantic and syntactic structure, forensic authorship, embedding space, and multi-modal (e.g., language-vision model) analysis.
@@ 9.6 Adversarial Robustness of TDA Features
Robustness to noise, particularly adversarial perturbations, has been an important research topic in NLP. While such robustness of TDA features is promising, there have been only a few works in this direction (Perez & Reinauer, 2022; Chauhan & Kaul, 2022). For instance,  Perez & Reinauer (2022) show that their topologically-augmented BERT model is much more robust than the base BERT model when tested against perturbations generated by TextAttack (Morris et al., 2020).  Chauhan & Kaul (2022) also shows that there are some weak correlations between persistent homology features of a trained BERT model and its adversarial robustness against several state-of-the-art attackers. However, all existing works only evaluate on BERT model with simple attack mechanism, missing other security scenarios such as poisoned or backdoor attacks. Therefore, we call for a more comprehensive robustness evaluation of TDA on NLP models, especially from the NLP, ML, and security communities.
@@ 9.7 Topological Deep Learning for NLP
Due to the benefits of TDA and deep learning, a new niche field is born - Topological Deep Learning (TDL) a “the collection of ideas and methods related to the use of topological concepts in deep learning” (Papamarkou et al., 2024). Initially, TDL is described as an ensemble of topological features extracted by TDA techniques such as persistent homology and deep learning features. In this setting, TDL is a traditional deep learning model with extra features (i.e., TDA-extracted features). However, as the field has advanced, a new definition for TDL has emerged - “the collection of ideas and methods related to the use of topological concepts in deep learning” (Papamarkou et al., 2024). TDL allows a deep learning model to be integrated more deeply with concepts of algebraic topology, such as the introduction of simplicial neural networks (NNs), which are NNs with layers made up of simplicial complexes. This deeper integration of TDA into NNs makes TDL particularly useful for the explosion of high-dimensional data. These high-dimensional data need better tools for processing as the current tools shrink the dimension, resulting in information loss. In NLP, one particular approach to integrate TDA with high-dimensional NLP embeddings has been the utilization of text in graphical forms, which have been shown to yield better results than directly using texts as a sequence of tokens (Zhong et al., 2020; Liu et al., 2023; Phan et al., 2023). Nevertheless, more research is still needed to validate such an approach.